OpenStudy (anonymous):

evaluate the log expression 3log(base 3) 21

OpenStudy (ghazi):

\[3\log _{3}21= 3(\frac{ \log 21 }{ \log 3 })=3(\frac{ \log(7*3) }{ \log3 })=3(\log(7)+\log(3))/\log 3\]

OpenStudy (ghazi):

\[\log _{b} A=\frac{ \log A }{ \log b }\] use this

OpenStudy (ghazi):

\[\log (AB)= \log A+ \log B\]

OpenStudy (anonymous):

of if you want a decimal approximation for this, you can compute \[3\times \frac{\ln(21)}{\ln(3)}\] by the change of base formula either way

OpenStudy (anonymous):

is it 21? because that is what I worked out.

OpenStudy (ghazi):

\[3(\frac{ \log 7 }{ \log3 }+\frac{ \log 3 }{ \log 3 })\] now you can do the rest

OpenStudy (anonymous):

if it is \[3\log_3(21)\] then no

OpenStudy (anonymous):

hmm then should it be 3??

OpenStudy (ghazi):

no

OpenStudy (anonymous):

would the answer happen to be in the negatitive form?

OpenStudy (ghazi):

no

OpenStudy (anonymous):

okay. so then I just got 1.04

OpenStudy (ghazi):

it's 5.38